The rate constant for this first‑order reaction is 0.850M ^−1 * s^-1 at 300 degrees Celsius.

A⟶products

How long, in seconds, would it take for the concentration of A to decrease from 0.710M to 0.290M?
1
Expert's answer
2020-02-27T09:02:10-0500

From First order rate law

T=(2.303/K) (log(a/(a-x)))

Where k=rate constant

T=time taken

a=initial concentration

a-x= reduced concentration

But for 1st order reaction the unit of the rate constant must be sec-1

Here the unit is M-1sec-1 which is the unit of 2nd order rate constant.

According to the question

T=(2.303/0.850)(log(0.710/0.290))

=1.053 sec required